Rate of formation of SO3 according to the reaction 2SO2+O2=2SO3 is 1.6×10-3 kg/min. What is the rate at which SO2 reacts?

Grade:12

2 Answers

Yash Donga
6 years ago
Verified answers contain reliable, trustworthy information vouched for by a hand-picked team of experts. Brainly has millions of high quality answers, all of them carefully moderated by our most trusted community members, but verified answers are the finest of the finest.
Artika
5 years ago
  • Rate of reaction = 1/2 rate of appearance of SO3 = 1.6×10-3
  • 1/2 Rate of dissapearance of SO2 = 1/2 Rate of appearance of SO3 = 1.6×10-3